OpenStudy (anonymous):

1. - 5(2x+2) - 10x - 17 (1 point)

OpenStudy (lilg132):

ok first expand the brackets -5 * 2x = -5 * 2 =

jhonyy9 (jhonyy9):

-10x-10-10x-17=-20x-27